A movie theater charges $5 admission for an adult and $3 for a child. If 700 tickets were sold and the total revenue received was $2900, how many child tickets were sold?

Answer by richwmiller(17219)

a+b=700,
5*a+3*b=2900
a=700-b
5*(700-b)+3*b=2900
3500-5b+3*b=2900
-2*b=-600
b=300
a=700-b
a=400 at $5 and b=300 at $3
check
400+300=700
5*a+3*b=2900
5*400+3*300=2900
2000+900=2900
2900=2900
ok